The EMF of an electric cell with
internal resistance of 2 Ω is 3.5V. If a
current of 0.5A flows when
a resistor is connected across it. Find the
resistance of the resistor.
1
Expert's answer
2020-12-06T17:21:41-0500

"I=\frac{E}{R+r}\to R=\frac{E-Ir}{I}=\frac{3.5-0.5\cdot 2}{0.5}=5(\Omega)"